OpenStudy (anonymous):

Part 1: Find the polynomial f(x) that has the roots of –2, 5 of multiplicity 2. (4 points) Part 2: Explain how you would verify the zeros of f(x). (4 points) I forgot how to do this help pls

OpenStudy (campbell_st):

well 1 solution is f(x) = (x + 2)(x - 5)^2 to prove distribute to get the polynomial and then use the factor theorem by finding f(-2) and f(5) if both give an answer of zero they are factors.

OpenStudy (anonymous):

oh then set them equal to 0 right ? okay thanks

OpenStudy (campbell_st):

no... when you have f(x) = x^3 +....... substitute x = -2 or f(-2) and show it is equal to zero.... this is known as the factor theorem... (x + a) is a factor of a polynomial if f(-a) = 0

OpenStudy (anonymous):

ohhhh okay its coming back to me, thanks so much :)